# Dissecting Lepton Number Violation in the Left-Right Symmetric Model: $0\nu\beta\beta$ decay, Møller scattering, and collider searches

### v1

@@ Abstract
In the context of the left-right symmetric model, we study the interplay of neutrinoless double beta ($0\nu\beta\beta$) decay, parity-violating Møller scattering, and high-energy colliders, resulting from the Yukawa interaction of the right-handed doubly-charged scalar to electrons, which could evade the severe constraints from charged lepton flavor violation. The ${0\nu\beta\beta}$ decay amplitude receives additional contributions from right-handed sterile neutrinos. The half-life, calculated in the effective field theory (EFT) framework, allows for an improved description of the contributions involving non-zero mixing between left- and right-handed $W$ bosons and those arising from exchanging a light right-handed neutrino. We find that the relative sensitivities between the low-energy (or high-precision) and high-energy experiments are affected by the left-right mixing. On the other hand, our results show how the interplay of collider and low-energy searches provides a manner to explore regions that are inaccessible to ${0\nu\beta\beta}$ decay experiments.
preprint: ACFI-T24-06, INT-PUB-24-040
I Introduction
@@
Explaining the origin of neutrino masses is a key open problem in particle and nuclear physics. If neutrinos are Majorana fermions, the mass term violates the lepton number by two units, $\Delta L=2$. The simplest tree-level realizations of the corresponding lepton-number-violating (LNV) Weinberg operator Weinberg (1979) at dimension-5 are the type-I Minkowski (1977); Gell-Mann et al. (1979); Yanagida (1979); Glashow (1980); Mohapatra and Senjanovic (1980), type-II Konetschny and Kummer (1977); Magg and Wetterich (1980); Schechter and Valle (1980); Mohapatra and Senjanovic (1981); Lazarides et al. (1981); Cheng and Li (1980), and type-III Foot et al. (1989); Ma (1998) seesaw mechanisms. Thus, in order to understand how neutrino masses are generated, it is essential to determine whether or not neutrinos are Majorana particles. The search for the neutrinoless double beta (${0\nu\beta\beta}$) decay process $(A,Z)\to(A,Z+2)+e^{-}+e^{-}$ Schechter and Valle (1982a) provides one of the most powerful experimental probes of this question.
@@
However, the detailed interpretation of ${0\nu\beta\beta}$ decay searches depends on the underlying mechanism. Beyond the “standard mechanism” involving the exchange of three light Majorana neutrinos, ${0\nu\beta\beta}$ decay may also receive non-standard contributions from other $\Delta L=2$ LNV sources beyond the Standard Model (BSM) if the associated mediator is sufficiently light. This possibility has been extensively studied in Refs. Hirsch et al. (1996); Faessler et al. (1997); Prezeau et al. (2003); Allanach et al. (2009); Tello et al. (2011); Bonnet et al. (2013); Helo et al. (2013a, b); Peng et al. (2016); Cirigliano et al. (2017, 2018, 2018); Li et al. (2021); Harz et al. (2021); Li et al. (2022a); Graesser et al. (2022); Li et al. (2022b); Bolton et al. (2022). A systematic description of ${0\nu\beta\beta}$ decay can be performed in the effective field theory (EFT) approach, which provides an end-to-end framework from the LNV scale to the nuclear scale Prezeau et al. (2003); Cirigliano et al. (2017, 2018).
@@
In order to distinguish different contributions to ${0\nu\beta\beta}$ decay, one needs to study the correlated signals in the high-energy and low-energy (or high-precision) frontiers Cirigliano et al. (2022). Complementary collider probes of LNV responsible for ${0\nu\beta\beta}$ decay at colliders have been explored in the context of ultraviolet (UV) complete models Prezeau et al. (2003); Allanach et al. (2009); Tello et al. (2011); Cirigliano et al. (2018); Li et al. (2021, 2022b); Bolton et al. (2022) and simplified models Bonnet et al. (2013); Helo et al. (2013a, b); Peng et al. (2016); Harz et al. (2021); Li et al. (2022a); Graesser et al. (2022). At the Large Hadron Collider (LHC), searching for same-sign charged leptons pairs plus di-jets Keung and Senjanovic (1983) provides a unique test of TeV-scale lepton number violation Helo et al. (2013b); Peng et al. (2016); Nemevšek et al. (2018); Harz et al. (2021); Nemevšek and Nesti (2023). Besides the LNV searches, the lepton-number-conserving (LNC) searches, including direct searches for new particles and precision tests of SM interactions, offer a valuable diagnostic of the BSM models, see for example Refs. Helo et al. (2013b); Dev et al. (2018); Du et al. (2022).
@@
Working within the framework of the left-right symmetric model (LRSM) Mohapatra and Pati (1975); Mohapatra and Senjanovic (1980, 1981), we will study ${0\nu\beta\beta}$ decay and its interplay with precision measurements of parity-violating Møller scattering and collider searches, focusing on the doubly-charged scalar in the right-handed sector. The LRSM provides a natural origin of neutrino masses in the type-I and type-II seesaw mechanisms, in which right-handed neutrinos and scalar triplets are introduced. Due to the right-handed charged currents, non-standard contributions to ${0\nu\beta\beta}$ decay arise Mohapatra and Senjanovic (1981), which may dominate over those from the exchange of light active neutrinos Prezeau et al. (2003); Tello et al. (2011); Li et al. (2021). The scalar potential of the model is provided with doubly-charged scalars from the left-right scalar triplets. The mass and Yukawa couplings of the doubly-charged scalar in the left-handed sector are severely constrained by charged lepton flavor violation (CLFV) Tello et al. (2011); Dinh et al. (2012); Barry and Rodejohann (2013); Dev et al. (2018); Barrie and Petcov (2023). On the contrary, doubly-charged scalar in the right-handed sector can evade such constraints if the left-right symmetry is not imposed in the Yukawa interactions Dev et al. (2018); Fukuyama et al. (2022) as compared to the minimal LRSM Mohapatra and Senjanovic (1980, 1981).
@@
The interaction of the doubly-charged scalar in the right-handed sector (hereafter right-handed doubly-charged scalar) with electrons not only induces a contribution to ${0\nu\beta\beta}$ decay but also to parity-violating asymmetries in polarized Møller scattering and the same-sign di-electron plus di-jet signature in proton-proton collisions. Thus, both observables provide complementary probes of the ${0\nu\beta\beta}$ decay mechanism in this context. The corresponding interplay among these observables has been discussed in Ref. Dev et al. (2018), where the analysis was focused on the regime where the doubly-charged scalar contribution dominated the ${0\nu\beta\beta}$ decay amplitude and no left-right mixing was considered. Here, we will investigate the connections with the following new highlights:
@@
  • A possible non-zero $W_{L}-W_{R}$ mixing (also named left-right mixing) is considered. It has been shown in the EFT framework Prezeau et al. (2003); Li et al. (2021); Cirigliano et al. (2018) that the contribution from the simultaneous exchange of $W_{L}$ and $W_{R}$ and right-handed neutrinos can dominate over the others to ${0\nu\beta\beta}$ decay because of a chiral enhancement at the hadronic level.
@@
  • As a consequence of using the EFT description in ${0\nu\beta\beta}$ decay, we correctly identify the parameter region where the right-handed neutrino transition between a light (dim-6) and a heavy (dim-9) EFT description exists, which proves to be essential for the sensitivity of ${0\nu\beta\beta}$-decay searches.
The remainder of the paper is arranged as follows. In Sec. II, we review the interactions, neutrino mixing, and left-right mixing in the LRSM. In Sec. III, we discuss the current constraints and future prospects at high-energy colliders. In Sec. IV, we discuss the current constraints and future prospects in low-energy experiments, including ${0\nu\beta\beta}$ decay and parity-violating Møller scattering. In Sec. V, the results of ${0\nu\beta\beta}$ decay and combined sensitivities are presented and discussed in detail. We present our conclusions in Sec. VI.
II Left-Right Symmetric Model
The LRSM Pati and Salam (1974); Mohapatra and Pati (1975); Senjanovic and Mohapatra (1975); Senjanovic (1979); Mohapatra and Senjanovic (1980, 1981) extends the SM gauge group to $SU(3)_{\rm C}\times SU(2)_{\rm L}\times SU(2)_{\rm R}\times U(1)_{B-L}$, where $B$ and $L$ denote the SM abelian baryon and lepton quantum numbers. Henceforth, we denote the field representation under the $SU(2)_{\rm R,L}$ and $U(1)_{B-L}$ groups by $(X_{\rm R},Y_{\rm L},Z_{B-L})$. The scalar sector consists of one bidoublet $\Phi\in(2,2^{\ast},0)$,
@@ II Left-Right Symmetric Model
[…]
Right-handed (RH) neutrinos appear in the doublet $L_{R}=(N,\ell_{R})^{\sf T}$ and are thus charged under $SU(2)_{\rm R}$ and couple to right-handed gauge bosons. It also couples to the scalar sector via the Yukawa interactions
@@
After the electroweak symmetry breaking (EWSB), the leptonic Yukawa interactions in Eq. (II) give rise to neutrino masses through the type-I Minkowski (1977); Gell-Mann et al. (1979); Yanagida (1979); Glashow (1980); Mohapatra and Senjanovic (1980) and type-II Konetschny and Kummer (1977); Magg and Wetterich (1980); Schechter and Valle (1980); Mohapatra and Senjanovic (1981); Lazarides et al. (1981); Cheng and Li (1980) seesaw mechanisms in terms of the symmetric $6\times 6$ matrix Schechter and Valle (1982b); Xing and Zhou (2011)
@@
where the Dirac and Majorana mass matrices are given by $M_{D}=(\kappa\mathbf{\Gamma_{l}}+\kappa^{\prime}\mathbf{\widetilde{\Gamma}_{l}})/\sqrt{2}$ and $M_{L,R}=\sqrt{2}\,\,\mathbf{f_{L,R}}\,v_{L,R}$, respectively. We diagonalize this neutrino mass matrix by using a $6\times 6$ unitary matrix $U$,
@@
The neutrinos $\nu_{1,2,3}$ are active neutrinos, while $\nu_{4,5,6}$ are sterile neutrinos. We can decompose the unitary matrix $U$ as Mitra et al. (2012)
@@
where $U_{\mathrm{PMNS}}$ is the usual $3\times 3$ neutrino-mixing matrix, i.e. Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix, $U_{R}$ is an additional $3\times 3$ matrix and the matrix $R=M_{D}M_{R}^{-1}$ gives rise to the active-sterile mixing. This particular parametrization allows us to explicitly show the seesaw mechanism by noticing that Mitra et al. (2012)
@@
We will assume the type-II seesaw dominance, that is the neutrino mass matrix
$M_{n}=\begin{pmatrix}M_{L}&0\\ 0&M_{R}\end{pmatrix}\,.$ (14)
From Eqs. (11), (12) and (13), we obtain
$U^{\dagger}_{\mathrm{PMNS}}M_{L}U_{\mathrm{PMNS}}^{*}$ $=\widehat{M}_{\nu}\,,$ $U^{\dagger}_{R}M_{R}U_{R}^{*}$ $=\widehat{M}_{N}\,.$ (15)
@@
where $\tan\zeta={\kappa\kappa^{\prime}}/{v_{R}^{2}}$. The charged gauge boson masses are
@@
where $g_{L,R}$ are the gauge couplings of $SU(2)_{\rm L,R}$, which are assumed to be equal $g\equiv g_{L}=g_{R}$. We introduce
@@
III Constraints and prospects at high-energy colliders
In this section, we study the current constraints and future sensitivities to the LRSM at high-energy colliders. Direct searches for the right-handed gauge boson $W_{R}$ at the LHC provide the most important diagnostic of the LRSM. The most stringent lower bound on the mass of $W_{R}$ is $6.4\,\mathrm{TeV}$ obtained by the ATLAS Collaboration Aad et al. (2023a) with the full Run-2 data of $138\,\text{fb}^{-1}$.
The same-sign dilepton searches at the LHC have excluded the doubly-charged scalar with mass below $1080\,\mathrm{GeV}$ using its decays into $\ell^{\pm}\ell^{\pm}$ (with $\ell=e,\mu$Aad et al. (2023b) with the full Run-2 data. A future high-energy $pp$ collider like the Super Proton-Proton Collider (SPPC) Tang et al. (2015) or the Future Circular Collider (FCC-hh) Abada et al. (2019), operating at a center-of-mass energy of $100\,\mathrm{TeV}$, is expected to probe the doubly-charged scalar to higher masses. Pair-production could occur via the Drell-Yan process, and the reach for the doubly-charged scalar mass with an integrated luminosity of 30 ab-1 could extend up to $5.1\,\mathrm{TeV}$ for $M_{W_{R}}=7\,\mathrm{TeV}$ and $2.5\,\mathrm{TeV}$ for $M_{W_{R}}\gtrsim 15\,\mathrm{TeV}$, respectively Dev et al. (2016).
In addition to the direct searches, the Bhabha scattering process $e^{+}e^{-}\to e^{+}e^{-}$ places an indirect constraint on the mass of the doubly-charged scalar. Ref. Nomura et al. (2018) obtained $M_{\delta_{R}^{\pm\pm}}/|(f_{R})_{ee}|>2.43\,\mathrm{TeV}$ based on the Large Electron-Positron (LEP) collider data Schael et al. (2013), improving the result derived in Ref. Dev et al. (2018). We follow the approach in Ref. Dev et al. (2018) to set a conservative limit for the reach of future $e^{+}e^{-}$ colliders by rescaling the LEP limit by a factor of

### v2

# Dissecting Lepton Number Violating Interactions in the Left-Right Symmetric Model: $0\nu\beta\beta$ decay, Møller scattering, and collider searches

@@ Abstract
In the context of the left-right symmetric model, we study the interplay of neutrinoless double beta ($0\nu\beta\beta$) decay, parity-violating Møller scattering, and high-energy colliders, resulting from the Yukawa interaction of the right-handed doubly-charged scalar to electrons, which could evade the severe constraints from charged lepton flavor violation. The $0\nu\beta\beta$} decay amplitude receives additional contributions from right-handed sterile neutrinos. The half-life, calculated in the effective field theory framework, allows for an improved description of the contributions involving non-zero mixing between left- and right-handed $W$ bosons and those arising from exchanging a light right-handed neutrino. We find that the relative sensitivities between the low-energy (or high-precision) and high-energy experiments are affected by the left-right mixing. On the other hand, our results show how the interplay of collider and low-energy searches provides a manner to explore regions that are inaccessible to $0\nu\beta\beta$} decay experiments.
1 Introduction
@@
Explaining the origin of neutrino masses is a key open problem in particle and nuclear physics. If neutrinos are Majorana fermions, the mass term violates the lepton number by two units, $\Delta L=2$. The simplest tree-level realizations of the corresponding lepton-number-violating (LNV) Weinberg operator [Weinberg:1979sa] at dimension-5 are the type-I [Minkowski:1977sc, Gell-Mann:1979vob, Yanagida:1979as, Glashow:1979nm, Mohapatra:1979ia], type-II [Konetschny:1977bn, Magg:1980ut, Schechter:1980gr, Mohapatra:1980yp, Lazarides:1980nt, Cheng:1980qt], and type-III [Foot:1988aq, Ma:1998dn] seesaw mechanisms. Thus, in order to understand how neutrino masses are generated, it is essential to determine whether or not neutrinos are Majorana particles. The search for the neutrinoless double beta ($0\nu\beta\beta$}) decay process $(A,Z)\to(A,Z+2)+e^{-}+e^{-}$ [Schechter:1981bd] provides one of the most powerful experimental probes of this question.
@@
However, the detailed interpretation of $0\nu\beta\beta$} decay searches depends on the underlying mechanism. Beyond the “standard mechanism” involving the exchange of three light Majorana neutrinos, $0\nu\beta\beta$} decay may also receive non-standard contributions from other $\Delta L=2$ LNV sources beyond the Standard Model (BSM) if the associated mediator is sufficiently light. This possibility has been extensively studied in Refs. [Hirsch:1996qw, Faessler:1996ph, Prezeau:2003xn, Allanach:2009iv, Tello:2010am, Bonnet:2012kh, Helo:2013dla, Helo:2013ika, Peng:2015haa, Cirigliano:2017djv, Cirigliano:2018yza, Cirigliano:2018yza, Li:2020flq, Harz:2021psp, Li:2021fvw, Graesser:2022nkv, Li:2022cuq, Bolton:2021hje]. A systematic description of $0\nu\beta\beta$} decay can be performed in the effective field theory (EFT) approach, which provides an end-to-end framework from the LNV scale to the nuclear scale [Prezeau:2003xn, Cirigliano:2017djv, Cirigliano:2018yza].
@@
In order to distinguish different contributions to $0\nu\beta\beta$} decay, one needs to study the correlated signals in the high-energy and low-energy (or high-precision) frontiers [Cirigliano:2022oqy]. Complementary collider probes of LNV responsible for $0\nu\beta\beta$} decay at colliders have been explored in the context of ultraviolet (UV) complete models [Prezeau:2003xn, Allanach:2009iv, Tello:2010am, Cirigliano:2018yza, Li:2020flq, Li:2022cuq, Bolton:2021hje] and simplified models [Bonnet:2012kh, Helo:2013dla, Helo:2013ika, Peng:2015haa, Harz:2021psp, Li:2021fvw, Graesser:2022nkv]. At the Large Hadron Collider (LHC), searching for same-sign charged leptons pairs plus di-jets [Keung:1983uu] provides a unique test of TeV-scale lepton number violation [Helo:2013ika, Peng:2015haa, Nemevsek:2018bbt, Harz:2021psp, Nemevsek:2023hwx]. Besides the LNV searches, the lepton-number-conserving (LNC) searches, including direct searches for new particles and precision tests of SM interactions, offer a valuable diagnostic of the BSM models, see for example Refs. [Helo:2013ika, Dev:2018sel, Du:2022vso].
@@
Working within the framework of the left-right symmetric model (LRSM) [Pati:1974yy, Mohapatra:1974gc, Mohapatra:1974hk, Senjanovic:1975rk, Senjanovic:1978ev], we will study $0\nu\beta\beta$} decay and its interplay with precision measurements of parity-violating Møller scattering and collider searches, focusing on the doubly-charged scalar in the right-handed sector. The LRSM provides a natural origin of neutrino masses in the type-I and type-II seesaw mechanisms, in which right-handed neutrinos and scalar triplets are introduced. Due to the right-handed charged currents, non-standard contributions to $0\nu\beta\beta$} decay arise [Mohapatra:1980yp], which may dominate over those from the exchange of light active neutrinos [Prezeau:2003xn, Tello:2010am, Li:2020flq]. The scalar potential of the model is provided with doubly-charged scalars from the left-right scalar triplets. The mass and Yukawa couplings of the doubly-charged scalar in the left-handed sector are severely constrained by charged lepton flavor violation (CLFV) [Tello:2010am, Dinh:2012bp, Barry:2013xxa, Dev:2018sel, Barrie:2022ake]. On the contrary, a doubly-charged scalar in the right-handed sector could have a mass much lower than its counterpart in the left-handed sector, as is the case in non-minimal LRSMs [Chang:1983fu, Sahu:2006pf, Deppisch:2014zta, Pritimita:2016fgr, Akhmedov:2024rvp] and its interactions can evade the CLFV constraints if the left-right symmetry is not imposed in the Yukawa interactions [Dev:2018sel, Akhmedov:2024rvp] [^1] In the minimal LRSM (mLRSM) [Mohapatra:1979ia, Mohapatra:1980yp], the CLFV constraints on the right-handed doubly-charged scalar can also be avoided by assuming that its Yukawa coupling is diagonal [Fukuyama:2022dhe, Belfkir:2023lot], and neutrinos obtain masses via the type-I seesaw mechanism. .
The interaction of the doubly-charged scalar in the right-handed sector (hereafter right-handed doubly-charged scalar) with electrons not only induces a contribution to $0\nu\beta\beta$} decay but also to parity-violating asymmetries in polarized Møller scattering, Bhabha scattering, and the same-sign di-electron plus di-jet signature in proton-proton collisions. If Møller experiments or future Bhabha scattering in $e^{+}e^{-}$ collisions observe deviations from the SM, such signals could be explained within non-manifest LRSM. To further determine whether the underlying physics is associated with lepton number violation, one can search for the $0\nu\beta\beta$} decay signals. Conversely, if a signal is observed in $0\nu\beta\beta$} decay, its origin in the LRSM can be tested by searching for deviations in Møller or Bhabha scattering experiments. In that case, compatibility with neutrino masses and CLFV constraints would again point toward some variant of the LRSM, possibly incorporating $D$-parity breaking [Chang:1983fu, Sahu:2006pf, Deppisch:2014zta, Akhmedov:2024rvp].
@@
The corresponding interplay among these observables has been discussed in Ref. [Dev:2018sel], where the analysis was focused on the regime where the doubly-charged scalar contribution dominated the $0\nu\beta\beta$} decay amplitude, and no left-right mixing was considered. Here, we will investigate the connections with the following new highlights:
@@
  • A possible non-zero $W_{L}-W_{R}$ mixing (also named left-right mixing) is considered. It has been shown in the chiral EFT framework [Prezeau:2003xn, Li:2020flq, Cirigliano:2018yza] that the contribution from the simultaneous exchange of $W_{L}$ and $W_{R}$ and right-handed neutrinos can dominate over the others to $0\nu\beta\beta$} decay because of a chiral enhancement at the hadronic level.
@@
  • As a consequence of using the EFT description in $0\nu\beta\beta$} decay, we identify the parameter region where the right-handed neutrino transition between a light (dim-6) and a heavy (dim-9) EFT description exists, which proves to be essential for the sensitivity of $0\nu\beta\beta$}-decay searches.
We obtain improved results in previously explored parameter regions and extend our exploration to the parameter space where the $0\nu\beta\beta$} decay amplitude is predominantly governed by light sterile neutrinos. This study also contributes to understanding the complementarity of $0\nu\beta\beta$} decay and low- and high-energy probes. On the one hand, we expand the relatively scarce literature connecting with PV Møller scattering, offering a more integrated treatment of these low-energy processes and providing important context for their interpretation. On the other hand, we update existing collider constraints from dedicated studies [Dev:2016dja, Nomura:2017abh, Dev:2018sel], although that was not the exclusive focus of this work. We carefully explain the different collider constraints shown in our key figures, which now incorporate the very light sterile neutrino region and the small Yukawa coupling region. These areas have not been thoroughly explored in the literature before.
The remainder of the paper is arranged as follows. In Sec. 2, we review the interactions, neutrino mixing, and left-right mixing in the LRSM. In Sec. 3, we discuss the current constraints and future prospects at high-energy colliders. In Sec. 4, we discuss the current constraints and future prospects in low-energy experiments, including $0\nu\beta\beta$} decay and parity-violating Møller scattering. In Sec. 5, the results of $0\nu\beta\beta$} decay and combined sensitivities are presented and discussed in detail. We present our conclusions in Sec. 6.
2 Left-Right Symmetric Model
The LRSM [Pati:1974yy, Mohapatra:1974gc, Mohapatra:1974hk, Senjanovic:1975rk, Senjanovic:1978ev] extends the SM gauge group to $SU(3)_{\rm C}\times SU(2)_{\rm L}\times SU(2)_{\rm R}\times U(1)_{B-L}$, where $B$ and $L$ denote the SM abelian baryon and lepton quantum numbers. Henceforth, we denote the field representation under the $SU(2)_{\rm R,L}$ and $U(1)_{B-L}$ groups by $(X_{\rm R},Y_{\rm L},Z_{B-L})$. The scalar sector consists of one bidoublet $\Phi\in(2,2^{\ast},0)$,
@@ 2 Left-Right Symmetric Model
where all parameters are real [Deshpande:1990ip]. The right-handed scalar triplet vev, $v_{R}$, is taken to lie well above the electroweak scale $v=246\,\mathrm{GeV}$. The bidoublet vevs, $\kappa$ and $\kappa^{\prime}$, satisfy $\sqrt{\kappa^{2}+\kappa^{\prime\,2}}=v$. The phases $\alpha$ and $\theta_{L}$ induce CP violation and will be set to zero ($\alpha,\,\theta_{L}=0$) for the purposes of this study. The scalar potential of the triplet and bidoublet sector follows the usual form [Deshpande:1990ip, Dekens:2015csa, Maiezza:2016bzp, Maiezza:2016ybz], leading to a spectrum that allows the doubly charged scalars to be relatively light [Akhmedov:2024rvp, Dev:2016dja]. This is because the squared mass, $M_{\delta_{R}^{\pm\pm}}^{2}\simeq\rho_{2}v_{R}^{2}$, is set by the dimensionless coupling $\rho_{2}$ from the scalar potential.
Right-handed neutrinos appear in the doublet $L_{R}=(N,\ell_{R})^{\sf T}$ and are thus charged under $SU(2)_{\rm R}$ and couple to right-handed gauge bosons. It also couples to the scalar sector via the Yukawa interactions
@@
After the electroweak symmetry breaking (EWSB), the leptonic Yukawa interactions in Eq. (2) give rise to neutrino masses through the type-I [Minkowski:1977sc, Gell-Mann:1979vob, Yanagida:1979as, Glashow:1979nm, Mohapatra:1979ia] and type-II [Konetschny:1977bn, Magg:1980ut, Schechter:1980gr, Mohapatra:1980yp, Lazarides:1980nt, Cheng:1980qt] seesaw mechanisms in terms of the symmetric $6\times 6$ matrix [Schechter:1981cv, Xing:2011zza]
@@
where the Dirac and Majorana mass matrices are given by $M_{D}=(\kappa\mathbf{\Gamma_{l}}+\kappa^{\prime}\mathbf{\widetilde{\Gamma}_{l}})/\sqrt{2}$ and $M_{L,R}=\sqrt{2}\,\,\mathbf{f_{L,R}}\,v_{L,R}$, respectively, [^2] If the left-right symmetry is not assumed in the Yukawa interactions, the phase $\theta_{L}$ contributes to $M_{L}$, but it does not affect the right-handed sector (unlike the scenario studied in Ref. [deVries:2022nyh]). . We diagonalize this neutrino mass matrix by using a $6\times 6$ unitary matrix $U$,
@@
The neutrinos $\nu_{1,2,3}$ are active neutrinos, while $\nu_{4,5,6}$ are sterile neutrinos. We can decompose the unitary matrix $U$ as [Mitra:2011qr]
@@
where $U_{\mathrm{PMNS}}$ is the usual $3\times 3$ neutrino-mixing matrix, i.e. Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix, $U_{R}$ is an additional $3\times 3$ matrix and the matrix $R=M_{D}M_{R}^{-1}$ gives rise to the active-sterile mixing. This particular parametrization allows us to explicitly show the seesaw mechanism by noticing that [Mitra:2011qr]
@@
[…]
@@
where $\tan\zeta={\kappa\kappa^{\prime}}/{v_{R}^{2}}$. Notice that, for an arbitrary CP-phase $\alpha$, one has $W_{L}^{\pm}=\cos\zeta\,W_{1}^{\pm}-\sin\zeta e^{\mp{\rm i}\alpha}\,W_{2}^{\pm}$ [deVries:2022nyh]. The charged gauge boson masses are
@@
where $g_{L,R}$ are the gauge couplings of $SU(2)_{\rm L,R}$. We introduce
@@
Notice that in the mLRSM, $g_{R}\equiv g_{L}$, while the LRSM with $D$-parity breaking predicts the relation $g_{R}=2g_{L}/3$ [Deppisch:2014zta]. In the context of an extended seesaw, however, there is no restriction on the relation between $g_{R}$ and $g_{L}$, as shown in Ref. [Pritimita:2016fgr]. For the purpose of our analysis, we will assume that $g_{R}=2g_{L}/3$. However, it is important to emphasize that our conclusions will remain valid regardless of the specific value chosen for $g_{R}$.
3 Constraints and prospects at high-energy colliders
In this section, we study the current constraints and future sensitivities to the LRSM at high-energy colliders. Direct searches for the right-handed gauge boson $W_{R}$ at the LHC provide the most important diagnostic of the LRSM. The most stringent lower bound on the mass of $W_{R}$ is $6.4\,\mathrm{TeV}$ obtained by the ATLAS Collaboration [ATLAS:2023cjo] with the full Run-2 data of $138\,\text{fb}^{-1}$.
The same-sign dilepton searches at the LHC have excluded the doubly-charged scalar with mass below $1080\,\mathrm{GeV}$ using its decays into $\ell^{\pm}\ell^{\pm}$ (with $\ell=e,\mu$[ATLAS:2022pbd] with the full Run-2 data. A future high-energy $pp$ collider like the Super Proton-Proton Collider (SPPC) [Tang:2015qga] or the Future Circular Collider (FCC-hh) [FCC:2018vvp], operating at a center-of-mass energy of $100\,\mathrm{TeV}$, is expected to probe the doubly-charged scalar to higher masses. Pair-production could occur via the Drell-Yan process, and the reach for the doubly-charged scalar mass with an integrated luminosity of 30 ab-1 could extend up to $5.1\,\mathrm{TeV}$ for $M_{W_{R}}=7\,\mathrm{TeV}$ and $2.5\,\mathrm{TeV}$ for $M_{W_{R}}\gtrsim 15\,\mathrm{TeV}$, respectively [Dev:2016dja].
In addition to the direct searches, the Bhabha scattering process $e^{+}e^{-}\to e^{+}e^{-}$ places an indirect constraint on the mass of the doubly-charged scalar. Ref. [Nomura:2017abh] obtained $M_{\delta_{R}^{\pm\pm}}/|(f_{R})_{ee}|>2.43\,\mathrm{TeV}$ based on the Large Electron-Positron (LEP) collider data [ALEPH:2013dgf], improving the result derived in Ref. [Dev:2018sel]. We follow the approach in Ref. [Dev:2018sel] to set a conservative limit for the reach of future $e^{+}e^{-}$ colliders by rescaling the LEP limit by a factor of